OpenStudy (anonymous):

KatrinaKaif how did u get 3t -7 = 2 - 2t?

OpenStudy (katrinakaif):

4t-7-t=7-2t-5 This is your orginal expression As you can see, there are 2 "t"s on the left side of the equal sign Collect the like terms first (on each side) Which means 4t - 7 - t = 3t - 7 Now let us move to the other side of the equal sign: 7 - 2t - 5 You can see that 5 can be substracted from 7 so : 7- 5 - 2t = 2 - 2t